OpenStudy (anonymous):

1.6 = 123.10 x S^0.54 x 0.151 How do you rearrange for S and find the answer, please show steps.

OpenStudy (anonymous):

@ganeshie8 @Geckoh2o @GameOver @HelpBlahBlahBlah @zscdragon

ganeshie8 (ganeshie8):

isolate S

ganeshie8 (ganeshie8):

Step1 : divide both sides by `123.10 x 0.151`

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

0.08607 = S^0.54

OpenStudy (anonymous):

@ganesjie8

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

take 1/0.54 power both sides

ganeshie8 (ganeshie8):

\(\large 0.08607 = S^{0.54}\) \(\large \left( 0.08607\right)^{\dfrac{1}{0.54}} = S\)